OpenStudy (anonymous):

Sorry, another question relating to the same thing. Just want to make sure I have the hang of it now! Solve the following systems of equations. Express your answer as an ordered pair with the format (a,b) with no spaces between the numbers or symbols. 3x+4y=16 -4x-3y=-19. I got the answer (4,1), is this correct? If not, could you help me out and point me in the direction of the right answer?

OpenStudy (anonymous):

For the first equation, 3x + 4y = 16: 3(4) + 4(1) ? 16 12 + 4 ? 16 16 = 16 Check. For the second equation, -4x - 3y = -19: -4(4) - 3(1) ? -19 -16 - 3 ? -19 -19 ? -19 Check. You are correct.